Grade 12th passPhysical Chemistry

Derive the expression relating the pH of buffer solutions for both acidic and basic buffer with the concentration of the components.

Dulal Pathak
7 Years agoGrade 12th pass

1 Answer

Ravleen Kaur
7 Years ago
According to the law of dissociation, the acid dissociation constantKacan be defined by the equation:

Ka= [H+][A–]/[HA]

where [H+] and [A–] are the concentrations of the ionized form of the acid, while [HA] is the concentration of the unionized form.

Taking log on both sides you get:

log Ka= log ([H+][A–]/[HA])

Splitting the log terms into separate components gives you:

log Ka= log [H+] + log ([A–]/[HA])

Since pKa= -log10Kaand pH = -log10[H+], therefore:

-pKa= – pH + log ([A–]/[HA]

Rearranging the terms we get the Henderson-Hasselbalch equation for acids as:

pH = pKa+ log ([A–]/[HA])

This equation can also be written as:

pH = pKa+ log [(salt)/(acid)]
or
pH = pKa+ log ([ionized]/[unionized])

Similarly, you can do for base.